# The Hitchhiker’s Guide to Program Analysis, Part II: Deep Thoughts by LLMs

### v2

@@ 3.2.1. Core Assumption and Rationale
This hypothesis enables SecIA to streamline analysis. By assuming the attacker achieves both arbitrary value control and can always reach the potential vulnerability, SecIA focuses solely on potential security impact. This permits early filtering of findings based purely on consequence, (safely) reducing subsequent analysis load. Critically, this approach defers the complex validation of actual program constraints—including path feasibility, value ranges, and importantly, whether those potentially protective checks or sanitizers can be manipulated through tainted data. This deferral mitigates false negatives (FNs) by preventing premature dismissal of vulnerabilities due to reliance on potentially bypassable checks or inaccurate LLM constraint reasoning about their effectiveness. The effectiveness of this design in reducing FNs is validated experimentally in §5.4.2.
For instance, in our motivating example, a simple analysis might deem the Tainted Loop Bound (TLB) on nums safe (i.e., not a bug) due to the goto statement potentially preventing an infinite loop condition. On the other hand, SecIA’s approach, guided by the AC-Hypo, intentionally ignores the effect of the goto statement at this stage (as its condition depends on the tainted data command_id) and proceeds to identify potential downstream effects (e.g., DoS) resulting from an arbitrarily large nums. The analysis of whether the goto check can indeed be bypassed under feasible program conditions is deferred.
3.2.2. Workflow
SecIA analyzes each location $K$ reported by the static analysis Taint Sink finding:
  • $\bullet$
    Forward Influence Analysis: Suppose the tainted data $v$ is hypothetically controlled by an attacker as $v_{atk}$, and the $Op(v)$ is a taint sink $K$ identified by static analysis. Starting from the result of $Op(v_{atk})$ at location $K$, trace how this tainted value propagates or influences subsequent program execution.
  • $\bullet$
    Identify Influenced Critical Operations: From the forward analysis, SecIA first pinpoints critical operations that are affected by the hypothetical attacker-controlled value $v_{atk}$. The analysis focuses on operations relevant to known vulnerability classes such as memory safety (e.g., arbitrary memory access) and DoS (e.g., insecure loop bounds).
  • $\bullet$
    Filter Benign Operations: SecIA leverages LLM-driven semantic understanding to distinguish and filter out operators that are actually benign. For instance, while the system flags potential risks when $v_{atk}$ impacts sensitive operations like array indexing (e.g., a[v]), it deliberately ignores cases recognized as safe, such as routine data structure traversals or searches (e.g., comparing element.id to v, a complete example appaers in §5.6). This selective filtering directs subsequent analysis (§3.3) to concentrate on issues that are more likely to present genuine security risks.
  • $\bullet$
    Output. The set of influenced Critical Operations identified in the last step. This set represents the downstream locations potentially impacted by the originally tainted data $v$ under the arbitrary control assumption.
In the Forward Influence Analysis step, the LLM is leveraged to identify potential influences, including indirect taint propagation. For example, in the motivating example, the LLM can recognize that the incremental pointer values is indirectly influenced by the tainted data nums through the GET_NEXT(). By doing so, SecIA can even identify unique bugs that are not directly related to the taint source (albeit we currently don’t have any such bugs in our experiments).
3.3. Constraint Assessor (ConA)
The Constraint Assessor (ConA) aims to identify whether the bug of COp can be triggered by analyzing the data constraints.
3.3.1. Design Rationale: Precondition and Postcondition of Constraints
As potential vulnerabilities often persist because the data constraints intended to prevent them are not truly effective in the specific context where the tainted data is used §3.1.2. Simply identifying the presence of a check or sanitization routine is insufficient. To rigorously evaluate effectiveness and guided by the paradigm of formal methods, the ConA is designed to determine:
  • $\bullet$
    Precondition. A constraint might exist in the code but be bypassed on a specific execution path leading to the potential vulnerability, or its activation might depend on configuration or state variables that are not set appropriately. Analyzing the Precondition, the conditions necessary for the constraint’s logic to execute addresses this.
  • $\bullet$
    Postcondition. Even if a constraint is activated, it might not be strict enough to prevent the specific value or range required to trigger the bug (e.g., checking x < 200 is ineffective if the vulnerability can happen if x > 100). Analyzing the Postcondition, the effect on the data’s possible values or state after the constraint operates addresses this.
Therefore, collecting and reasoning about both the Preconditions and Postconditions of constraints along the taint flow is essential for ConA to move beyond superficial checks and accurately assess whether the existing safeguards collectively guarantee the elimination of the vulnerability at the critical operation. This detailed steps is described in the following sections.
3.3.2. Overview of Workflow
As shown in Figure 3, the Constrant Assessor involves a four-step, LLM-guided workflow:
Figure 3. The workflow of Constraint Assessor (ConA)
@@ 3.3.2. Overview of Workflow
    Step 2: Backward Constraint Collection. Next, the analysis traces the tainted data flow path(s) backward from the COp towards the data’s source. Along this path, it identifies code segments—such as conditional statements, assertions, or calls to validation functions—that appear intended to act as constraints on the tainted data’s value or range before it is used at the COp.
@@
    Step 3: Constraint Effect Summarization. Each potential constraint identified in Step 2 is then analyzed in detail. This step aims to understand the constraint’s specific effect: Under which conditions (its activation precondition) does the constraint apply, and what impact does it have on the tainted variable’s possible numerical range (its range effect postcondition)?
@@
    Step 4: Final Vulnerability Evaluation. Finally, it performs a reasoned evaluation to determine if the vulnerability can be triggered. If still triggerable, the finding is classified as a “Potential Vulnerability.” Otherwise, if the LLM determines that the constraints effectively prevent the vulnerability from being triggered, the finding is classified as “Eliminated.”
This workflow leverages LLMs to interpret code, identify relevant patterns and constraints, and perform the heuristic reasoning required for each step. The following subsections (§3.3.3 through §3.3.6) elaborate on the specific design, rationale, heuristic considerations, and soundness implications inherent in each stage of this analysis process.
3.3.3. Step 1: Critical Operation Reachability Analysis.
@@
The LLM is then tasked (in Step 4) to evaluate these semantic constraints to assess potential impact. The natural language representation allows the LLM to leverage its understanding of the code’s intent and context, but introduces a higher degree of heuristics and potential uncertainty into the final evaluation, which must be acknowledged.
3.3.4. Step 2: Backward Constraint Collection.
@@
Leveraging Code Context and Structure: The LLM is tasked with identifying validation and sanitization operations of the tainted data. Specifically, the LLM looks for:
  • $\bullet$
    Conditional branches (if, switch) whose conditions involve the tainted variable or related data.
  • $\bullet$
    Assignments or operations that transform the tainted data.
  • $\bullet$
    Function calls that take the tainted data (or related data) as arguments.
Heuristic Identification in Domain Context: Unlike formal symbolic execution, which aims to exhaustively collect all mathematically precise path conditions, this LLM-based collection is heuristic. It may not identify every constraint on all possible paths. However, we leverage the LLM’s pattern recognition capabilities, particularly within our target domain (Linux kernel). The prompt encourages the LLM to pay attention to common kernel patterns, naming conventions (e.g., functions named check_..., validate_...), and comments that often signify sanitization or validation routines, even if they are complex to model formally. This allows the LLM to identify likely and semantically significant constraints that purely syntactic approaches might suffer from path explosion.
@@
Example: (Figure 3) Tracing back from critical_op(x) in caller(), the LLM identifies the call check_x(...) to be a potential constraint on x. It would then query the Information Agent for the definition of check_x() to analyze its body (as detailed in Step 3)
3.3.5. Step 3: Constraint Effect Summarization
@@
After collecting potential constraining code segments (like the function check_x) in Step 2, this step analyzes the effect of these segments on the tainted variable. The goal is to understand how different execution paths within these segments modify the possible range of the tainted variable, and under which conditions (preconditions) those paths are taken.
Traditional formal methods, like symbolic execution combined with abstract interpretation, analyze such segments rigorously. However, to ensure soundness and termination, particularly with loops or complex arithmetic, they often employ over-approximations for numerical ranges. While guaranteeing soundness (no false negatives), this necessary conservatism can sometimes limit precision, failing to prove tight constraints that actually hold, thus potentially leading to false positives later.
BugLens utilizes the LLM. We prompt the LLM to interpret the control flow and data transformations within the identified constraint segment to heuristically estimate the range effects. The prompt asks the LLM to first identify all major execution paths through the provided code segment (e.g., the body of check_x), and then:
  • $\bullet$
    For each path, determine the Path Condition (Precondition): The conditions on the parameters or relevant program state required to take the path.
  • $\bullet$
    For each path, determine the Estimated Effect on Tainted Data (Postcondition): Analyze operations like assertions (assert(x > 0)), assignments (x = sanitize(x)), and comparisons to infer the resulting constraint on the tainted variable’s range.
Table 1. The precondition and postcondition pairs for the check_x function. The oc and c stand for other_config and config respectively.
Path Precondition Postcondition
$\pi_{1}$ oc = SKIP $x\in(-\infty,+\infty)$
$\pi_{2}$ oc $\neq$ SKIP $\land$ c = 1 $x\in(0,100)$
$\pi_{3}$ oc $\neq$ SKIP $\land$ c = 2 $x\in(0,+\infty)$
$\pi_{4}$ oc $\neq$ SKIP $\land$ c $\notin\{1,2\}$ $x\in(-\infty,+\infty)$
@@
Example: (Figure 3) For the function check_x in our example, the goal is for the LLM to output a summary similar to Table 1, which maps path conditions (preconditions) to range effects (postconditions). The path $\pi_{1}$ and $\pi_{4}$ are bypass paths, which means that the check_x function will not effectively limit the range of x.
Heuristic Nature and Soundness Trade-offs: This step relies on the LLM’s code interpretation capabilities and is heuristic, particularly in its estimation of range effects and its handling of complex control flow. Similarly to the Step 1, these summaries could also be represented in natural language, which brings more flexibility but also more uncertainty.
Rationale: We employ this LLM-based heuristic approach acknowledging the trade-offs. It aims to overcome the potential precision limitations of conservative formal abstractions and achieve broader applicability in the complex Linux kernel domain, where precise, scalable formal analysis is challenging. The goal is significantly improved precision (fewer FPs) compared to baseline static analysis, accepting a carefully managed risk of heuristic errors (potential FNs). The practical rate of False Negatives is thoroughly examined in our experimental evaluation in §5.3 and §5.5.
3.3.6. Step 4: Final Vulnerability Evaluation.
@@
This final step synthesizes the findings from the previous analyses to determine if the identified constraints effectively neutralize the potential vulnerability associated with the Critical Operation (COp). The goal is to classify the initial finding as either “Eliminated” (constraints provably prevent the the vulnerability happening) or ”Potential Vulnerability” (no such guarantee found).
Specifically, the LLM is guided to perform the following assessment:
  • $\bullet$
    Analyze Constraint Activation: The first step is to determine if the precondition (after considering user/kernel control effects) is implied by the known COp’s Reachability Conditions. If it is, the constraint is active as an effective constraint. Otherwise, we need to further analyze its activation based on the following heuristic.
  • $\bullet$
    Heuristic to Untainted Data: For untainted data, use the heuristics (e.g., variable name, __user flag) to determine if the variable can be controlled by user, then use the worst-case assumption that the user can manipulate these variables and analyze if the precondition can be unsatisfied. And a possibility of unsatisfied precondition makes the constraint disqualified.
  • $\bullet$
    Analyze Constraint Effectiveness: Next is to determine if the postcondition of the remaining active constraints
    prevents the Bug Condition of the COp from ever being satisfied. For example, if the vulnerability happens when idx < 0 and the constraint’s effect is estimated as idx in $[0,100]$, the vulnerability is prevented.
Soundness Implications: This final evaluation relies on the LLM’s ability to perform heuristic logical reasoning. As a result, this step may introduce the following concerns: (1) The heuristic to untainted data (worst-case assumption for user control) may underestimate/overestimate constraint effectiveness, (2) errors in the LLM’s logical reasoning, range comparisons, or handling of complex conditions could lead to classification errors, and (3) it inherits any potential unsoundness from Step 3 caused by incorrect range estimations or missed bypass conditions .
@@ 3.4. Structured Analysis Guidance (SAG)
To enable the LLM to effectively perform the code analysis tasks in ConA, we employ Structured Analysis Guidance (SAG) — a prompting strategy with structured reasoning templates and few-shot examples. SAG provides: (1) Guided Stepwise Vulnerability Analysis with step-by-step instructions that decompose complex analysis tasks following established formal method principles (the analysis of precondition and postcondition, as described in §3.3), and (2) Guided Path Condition and Data Constraint Analysis demonstrating how to analyze challenging data constraints and path conditions from code, as described below.
@@ 3.4.1. Guided Path Condition Analysis
In the analysis of the path condition in Step 1 and Step 3 of ConA3.3.3 and §3.3.5), we guide the LLM using prompts designed to extract path conditions from the source code surrounding the operation of interest. For example, consider the path leading to the op(x) call within the caller function:
@@ 3.4.2. Guided Data Constraint Analysis
In the analysis of data constraints in Step 2 and Step 3 of ConA3.3.4 and §3.3.5), We consider the following data constraints:
@@
3.5. LLM Agent for Codebase Information
Following recent work (llift, ), we adopt an agentic design where the LLM acts as an interactive code analyst. It begins with a single function from the repository, identifies potential issues, and explicitly requests more context (e.g., other function definitions, struct layouts, global variables) if needed. The system then retrieves the relevant code snippet(s), and this iterative loop continues until the LLM has sufficient context to complete the analysis.
We provide a set of request types (function definitions, struct layouts, global variables, etc.), enabling the LLM to gather broader codebase information at any point. This design is both flexible and extensible: new request types can be supported by adding corresponding backend callbacks.
To facilitate these interactions, the knowledge retrieval system parses requests, locates relevant code, and formats responses for subsequent reasoning steps. We implement this system (about 500 LOC in Python) on top of CodeQuery (kopathy_ruben2020codequery_2025, ), extending its default functionality with custom handlers for specialized queries.
@@ 5.2. Experimental Setup
5.2.1. Evaluation Dataset: Full Kernel Analysis with static analyzers
Our study utilizes the Android kernel (Linux version 4.14.150, Google Pixel 4XL), which served as the testbed in the original Suture study (zhang_statically_2021, ). For our initial analysis, potentially informing RQ1 regarding baseline performance and the challenges of automated bug detection in this complex environment, we applied prior static analysis tools, Suture and CodeQL-SOD, to the Linux kernel device drivers. The key findings from this analysis provide important context:
@@ 5.2.1. Evaluation Dataset: Full Kernel Analysis with static analyzers
    CodeQL-SOD: We leverage CodeQL (codeql, ) as a complementary static analysis tool beyond Suture. Based on Backhouse et al.’s approach (backhouse_stack_2018, ), we implemented a simple inter-procedural taint tracking analysis (CodeQL-SOD) that traces data flows from ioctl entry points to copy_from_user sink functions to detect potential stack overflows. Running CodeQL-SOD on the Linux device drivers, it yields 11 potential bug reports. Our manual analysis confirmed 1 of these as a true positive (consistent with the known stack exhaustion bug reported in (backhouse_stack_2018, )). This corresponds to an FP rate of 90.9%.
5.2.2. Cost.
@@ 5.3. RQ1: Effectiveness
Table 2. Performance of BugLens on top of Suture and CodeQL-SOD.
@@
CodeQL-SOD 1 0 10 0 0.09 1.00 0.17
CodeQL-SODBugLens 1 9 1 0 0.50 1.00 0.67
5.3.1. Precision and Recall.
Table 2 shows the performance for the evaluated two static analyzers and our post-refinement method, BugLens. The results show BugLens significantly enhances precision of both Suture (0.10) and CodeQL-SOD (0.09). For CodeQL-SOD, the refinement (CodeQL-SODBugLens) increases precision substantially to 0.50 while not missing any real bugs detected before. For Suture, the precision is increased to 0.72 due to a drastic reduction in false positives (from 227 to 9). This refinement does not miss any exisiting bugs.
Moreover, noting that the semi-automated method in Suture, noted as SutureRP, actually shows a lower recall (0.83) than the BugLens refinement (1.0). This is because after examining the positive results of SutureBugLens, we found 4 cases of real bugs that were incorrectly classified as false positives during the manual inspection process in SutureRP. The results for SutureRP are estimated on the performance reported from its paper (zhang_statically_2021, ), as the semi-automated approach involves manual analysis steps that we did not replicate.
5.3.2. New Bugs.
As mentioned in the previous part, we find 4 more cases that are actually real bugs, which are previously classified as false positive by human inspection in Suture. Two bugs are from the sound subsystem, reported to the maintainers, while waiting for their feedback. One of them involves a data constraint that appears to sanitize a tainted value but can be bypassed due to subtle control-flow logic.

### v3

@@ 4.1.1. Core Assumption and Rationale
This hypothesis enables SecIA to streamline analysis. SecIA focuses solely on potential security impact. This permits early filtering of findings based purely on the consequence, (safely) reducing subsequent analysis load. Critically, this approach defers the complex analysis of actual program constraints—including path feasibility, value ranges, and importantly, whether those protective checks or sanitizers really work. This deferral mitigates false negatives (FNs) by preventing premature dismissal of vulnerabilities due to reliance on potentially bypassable checks or inaccurate LLM constraint reasoning about their effectiveness. The effectiveness is shown in §6.4.2.
4.1.2. Security Impact Analysis
SecIA performs a forward influence analysis to identify the influenced critical operations (COp) that are potentially affected by the tainted data $v$ from the sink location $K$. It then filters out the benign operations that are not security-sensitive based on whether this operation could (potentially) result in memory bugs (e.g., out-of-bound access, arbitrary memory access) or DoS. For instance, the jiffies case in §3.1.1 will be attributed as “not a bug”.
Figure 2. The workflow of Constraint Assessor (ConA) with an example, where the x is tainted and op(x) is the critical operation (COP). The sanitization function check_x(..., x) is affected by the config (noted as c) and other_config (noted as oc).
@@ 4.1.2. Security Impact Analysis
4.2. Constraint Assessor (ConA)
The Constraint Assessor (ConA) aims to identify whether the bug of COp can be triggered by analyzing the data constraints. As shown in Figure 2, the Constraint Assessor involves a four-step workflow:
@@ 4.2. Constraint Assessor (ConA)
    Step 2: Constraint Collection. Next, the analysis traces the tainted data flow path(s) backward from the COp towards the data’s source. Along this path, it identifies code segments—such as conditional statements, assertions, or calls to validation functions—that appear intended to act as constraints on the tainted data’s value or range before it is used at the COp.
@@
    Step 3: Constraint Effect Analysis. Each potential constraint identified in Step 2 is then analyzed in detail. This step aims to understand the constraint’s specific effect: Under what conditions (precondition) does the constraint need to satisfy, and what impact does it have on the tainted variable’s possible numerical range (postcondition)?
@@
    Step 4: Final Vulnerability Evaluation. Finally, it performs an evaluation to determine if the potential vulnerability can still be triggered. If yes, the case is classified as a “Potential Vulnerability.” Otherwise, if the LLM determines that the constraints effectively prevent the vulnerability from being triggered, the case is classified as “Eliminated.”
This workflow leverages LLMs to interpret code (mimicking a formal reasoning process), identify relevant patterns and constraints, and perform the reasoning required for each step.
We employ this LLM-based approach acknowledging the trade-offs of soundness and precision. ConA aims to overcome the potential precision limitations of conservative abstractions and achieve broader applicability in the complex Linux kernel. In §6.3 and §6.5, we thoroughly examine how our design in relaxing the strict over-approximation introduces only a reasonably small number of false negatives. In the following sections (§4.2.1 through §4.2.4), we elaborate on the specific design, rationale, heuristic considerations, and soundness implications inherent in each stage of this analysis process.
4.2.1. Step 1: Critical Operation Reachability Analysis
@@
The LLM is later asked (in Step 4) to evaluate these semantic constraints to assess the potential impact. The natural language representation allows the LLM to leverage its understanding of the code’s intent and context but introduces a higher degree of uncertainty in the final evaluation, which we must acknowledge.
4.2.2. Step 2: Backward Constraint Collection
@@
Example: In Figure 2, tracing back from critical_op(x) in caller(), the LLM identifies the call check_x(...) to be a potential constraint on x. It would query the PKA for the function definition and analyze its effect, as detailed in Step 3.
4.2.3. Step 3: Constraint Effect Analysis
@@
After collecting potential constraining code segments (like the function check_x()) in Step 2, this step analyzes the effect of these segments on the tainted variable. The goal is to understand how different execution paths within these segments modify the possible range of the tainted variable and under which conditions (preconditions) those paths are taken.
This step prompts the LLM to first identify all major execution paths through the provided code segment (e.g., check_x()), and then considers the preconditions and postconditions (data constraints of the tainted data) of each path.
@@
Example: In Figure 2, we ask the LLM to analyze the function check_x() to summarize its effects on the tainted variable x. Particularly, the path $\pi_{1}$ and $\pi_{4}$ are bypass paths, and check_x() will not effectively limit the range of x with these preconditions.
4.2.4. Step 4: Final Vulnerability Evaluation
@@
This final step synthesizes the precondition and postcondition from the previous analyses to determine if the identified constraints effectively neutralize the potential vulnerability associated with the Critical Operation (COp). And then it classifies the reported vulnerability as either “Eliminated” (constraints effectively prevent the vulnerability happening) or “Potential Vulnerability” (no effective constraints).
[…]
@@ 4.3. Structured Analysis Guidance (SAG)
To support the fine-grained code reasoning steps in ConA, we employ Structured Analysis Guidance (SAG) to scale the test-time compute of LLMs with structured reasoning templates and few-shot examples. At a high level, SAG grounds the reasoning procedures of LLMs with typical program analysis steps by eliciting more reasoning tokens during inference. Specifically, SAG assists Constraint Assessor (ConA) with the following two types of analysis: (i) Guided Stepwise Vulnerability Analysis with step-by-step instructions that decompose the analysis of precondition and postcondition (as described in §4.2), and (ii) Guided Path Condition and Data Constraint Analysis to demonstrate how to analyze challenging data constraints and path conditions from code.
@@ 4.3.1. Guided Path Condition Analysis
In the analysis of the path condition in Step 1 and Step 3 of ConA4.2.1 and §4.2.3), we guide the LLM using prompts designed to extract path conditions from the source code surrounding the operation of interest. For example, consider the path leading to the op(x) call within the caller function:
@@ 4.3.2. Guided Data Constraint Analysis
In the analysis of data constraints in Step 2 and Step 3 of ConA4.2.2 and §4.2.3), We consider the following data constraints:
@@
Formal program analysis naturally distinguishes between validation and sanitization; however, LLMs often confuse these concepts (e.g., missing transferred validation, treating sanitization as validation). SAG emphasizes these differences with few-shot examples, and therefore improves the LLMs’ reasoning on data constraints.
4.4. Project Knowledge Agent (PKA)
BugLens leverages Project Knowledge Agent (PKA), a simple LLM code agent equipped with custom tools to automatically navigate the codebase and retrieve the related code context during the analysis. It starts the analysis with the context of the sink and explicitly requests more context (e.g., function definitions get_func_def, global variable definition get_var_def, struct layouts) if needed. PKA then retrieves the relevant code snippets iteratively until the model self-decides that it has sufficient context to complete the analysis.
We provide a set of request types, enabling the LLM to gather broader codebase information at any point. This design is both flexible and extensible: new request types can be supported by adding corresponding backend callbacks. PKA searches the relevant code snippets based on CodeQuery (kopathy_ruben2020codequery_2025, ). We implement PKA in 500 LOC of Python code.
@@ 6.2. Experimental Setup
6.2.1. Evaluation Dataset: Linux Kernel Driver Analysis with static analyzers
Our study utilizes the Android kernel (Linux version 4.14.150, Google Pixel 4XL), which served as the testbed in the original Suture study (zhang_statically_2021, ). For our initial analysis, potentially informing RQ1 regarding baseline performance and the challenges of automated bug detection in this complex environment, we applied prior static analysis tools, Suture and CodeQL-OOB, to the Linux kernel device drivers. The key findings from this analysis provide important context:
@@ 6.2.1. Evaluation Dataset: Linux Kernel Driver Analysis with static analyzers
    CodeQL-OOB: We leverage CodeQL (codeql, ) as a complementary static analysis tool beyond Suture. Based on Backhouse et al.’s approach (backhouse_stack_2018, ), we implemented a simple inter-procedural taint tracking analysis (CodeQL-OOB) that traces data flows from ioctl entry points to pointer dereference and copy_from_user (two typical cases of OOB bugs) to detect potential stack overflows. Running CodeQL-OOB on the Linux device drivers, it yields 24 potential bug reports. Our manual analysis confirmed 1 of these as a true positive (consistent with the known stack overflow bug reported in (backhouse_stack_2018, )). This corresponds to an FP rate of 95.8%.
6.2.2. Cost
@@ 6.3. RQ1: Effectiveness
Table 1. Performance of BugLens on top of Suture and CodeQL-OOB.
@@
CodeQL-OOB 1 0 23 0 0.04 - -
CodeQL-OOBBugLens 1 22 2 0 0.33 1.00 0.5
6.3.1. Precision and Recall
Table 1 shows the performance for the evaluated two static analyzers and our post-refinement method, BugLens. The results show BugLens significantly enhances precision of both Suture (0.10) and CodeQL-OOB (0.04). For CodeQL-OOB, BugLens increases precision substantially to 0.33 while not missing any real bugs detected before. For Suture, the precision is increased to 0.72 due to a drastic reduction in false positives (from 227 to 9). This refinement does not miss any exisiting bugs.
Moreover, noting that the semi-automated method in Suture, noted as SutureRP, actually shows a lower recall (0.83) than the BugLens refinement (1.0). This is because after examining the positive results of SutureBugLens, we found 4 cases of real bugs that were incorrectly classified as false positives during the manual inspection process in SutureRP.
6.3.2. New Bugs
As mentioned earlier, we found 4 more cases that are real bugs, which previously classified as false positive by human inspection in Suture. Two bugs are from the sound subsystem, reported to the maintainers, while waiting for their feedback. One of them involves a data constraint that appears to sanitize a tainted value but can be bypassed due to subtle control-flow logic.